P=12 units:Bottom (x-1),side 5,side x
This is one of the harder eqautions for me.
Help would not only be appreciated,but desired
Thank you:)

Answer by ankor@dixie-net.com(12692)

P=12 units:Bottom (x-1),side 5,side x
:
I understand this to mean the perimeter of a triangle, with the given sides = 12
:
The perimeter is the sum of all 3 sides:
(x-1) + 5 + x = 12
:
Add the x's,
x + x - 1 + 5 = 12
:
2x + 4 = 12
:
Subtract 4 from both sides:
2x = 12 - 4
:
2x = 8
:
Divide both sides by 2 and you have:
x = 8/2
:
x = 4
:
:
Check our solution, substitute 4 for x: (4-1) + 5 + 4 = 12